# The AdEMAMix Optimizer: Better, Faster, Older

### v1

@@ 1 Introduction
The use of momentum in optimization is grounded in the varying nature of gradients. As local linear approximations of the loss landscape, their information can quickly become outdated as the optimization process progresses (Pascanu et al., 2013). For this reason, practitioners typically employ moderate momentum values (i.e. $\beta\approx 0.8$ or $0.9$), effectively creating a moving average of recent gradients while discarding older information. Selecting larger $\beta$ values seems counter-intuitive, as it would suggest that older gradients maintain their relevance over extended periods of training. While it is tempting to see the use of small $\beta$s as a confirmation of the limited temporal relevance of gradients, our work reveals instead that older gradients can efficiently be used. When we increase $\beta$, we decrease the relative importance of recent gradients, and the iterate now fails to respond to local changes in the loss landscape. We observe that a single EMA cannot both give a significant weight to recent gradients, and give a non-negligible weight to older gradients (see Fig. 3(a)). However, a linear combination between a “fast-changing” (e.g. $\beta=0.9$) and a “slow-changing” (e.g. $\beta=0.9999$) EMA allows the iterate to beneficiate from (i) the great speedup provided by the larger (slow-changing) momentum, while (ii) still being reactive to small changes in the loss landscape (fast-changing). More precisely, we find the following statement to convey an important intuition behind this approach: {adjustwidth}2em2em While changing the direction of the slow momentum is difficult, any adjustment orthogonal to that direction is easy—which favors fast progress in sinuous canyon-like landscapes. A toy illustration of this can be seen in Fig. 2. Based on this idea, we propose AdEMAMix (Adaptive EMA Mixture), a novel Adam based optimizer which successfully leverages very old gradients to reach better solutions.
@@
  • We show AdEMAMix forgets the training data at a slower pace when compared to Adam (see Fig. 4).
@@ 2 Related Work
Works incorporating an additional momentum term. Lee et al. (2024) introduce Grokfast, which uses a pre-filtering step on the gradient to amplify the low frequencies and help solve groking. When combined with Adam, it effectively applies the Adam’s EMAs on top of another gradient averaging method (e.g. EMA for Grokfast-EMA). Somewhat similarly, Chen et al. (2023b) refer to the Double EMA (DEMA) used in some finance applications (Mulloy, 1994) as one motivation for their AdMeta optimizer. Our motivation behind AdEMAMix is to combine both a high sensitivity to the recent gradients as well as incorporating very distant gradient, in this respect, using nested EMAs is not the right candidate as it reduces the influence of recent gradients. A more detailed review of AdMeta and DEMA is in App. C.3.1.
@@ 3 Our method: AdEMAMix
Figure 3: Limitation of EMAs, constant $\eta$-scheduler, & Mamba results. In (a), we plot the weights $w_{t}$—for each past gradient $g_{t}$—given by different EMAs after $10k$ steps. For a given $\beta$, $\text{EMA}(\beta,{\bm{g}}^{(0)},\ldots,{\bm{g}}^{(T)})=\sum_{i=0}^{T}\beta^{i}(1-\beta){\bm{g}}^{(T-i)}$, which decays the contribution of past gradients exponentially. A small $\beta$ (e.g. $0.9$) will give a high weight to the immediate past and negligible weights to older timesteps. In contrast, a high $\beta$ (e.g. $0.9999$) is giving a relatively uniform, yet non-negligible weight to all gradients. No $\beta$ value can simultaneously give a high weight to the immediate past and a non-negligible weight to very old timesteps. In (b), we train multiple $1.3$B language models using $3k$ steps of warmup and then a constant learning rate $\eta=10^{-4}$. This allows us to observe the gap between AdamW and AdEMAMix without the cosine decay as a confounder. We still use schedulers for $\alpha$ and $\beta_{3}$ with $T_{\alpha,\beta_{3}}=500k$, $\alpha=5$. Similar to Zhai et al. (2022); Hu et al. (2024); Hägele et al. (2024), we decay the learning rate linearly at $t=1$M and $t=1.3$M. The loss-gap between AdamW and AdEMAMix increases at first, and then remains constant. AdEMAMix still outperforms AdamW after decaying the learning rate. See App.C.1.7 to see the impact of the linear decay duration. In (c), we train $168$M parameter Mamba models, showing how AdEMAMix’s performances can generalize outside of the Transformer architecture.
@@
Tackling early training instabilities. Early training instabilities are commonplace when training deep models, and empirically motivated the introduction of methods such as learning rate warmup and gradient clipping. Gilmer et al. (2022) show how the use of learning rate warmup can be justified from a curvature perspective; allowing the iterates to move to parts of the optimization landscape where larger learning rates are stable. While we use learning rate warmup in all our experiments, we still noticed AdEMAMix models using a large $\beta_{3}$ would diverge early. This, despite not using bias correction over ${\bm{m}}_{2}$, which lets the momentum buffer fill itself slowly. Those failed runs are characterized by updates of large magnitudes in the early phase of training (see App. C.1.7, Fig. 22). For this reason, we progressively increase the values of $\beta_{3}$ and $\alpha$ using schedulers. For $\alpha$ we use a linear scheduler. A linear scheduler for $\beta_{3}$ would be ill-fitted as the same increment of $\beta_{3}$ have a different impact for different values of $\beta_{3}$. For instance, observe that an increase of $\beta$ of $\delta_{\beta}=0.0001$ barely increases the $t_{half}$ for $\beta=0.9$, while $0.999\rightarrow 0.999+\delta_{\beta}$ increases the $t_{half}$ of $77$. For this reason, we design the $\beta_{3}$ scheduler to increase $t_{half}$ linearly (see App. A.1 for a derivation of that scheduler). The two schedulers are summarized below:
@@
Computational overheads. Adding an additional EMA requires additional memory and compute. The added compute is negligible in comparison to what is required for the forward-backward, and has little impact over the total runtime as shown in Fig. 5(a). Moreover—when considering larger distributed setups—it is worth noting that AdEMAMix is not increasing communication (gradient reduction) over Adam. Therefore, we expect the runtime overhead of AdEMAMix to shrink in those settings, as data movements occupy a larger fraction of the total runtime. A more significant overhead is in terms of memory, as AdEMAMix requires to allocate ${\bm{m}}_{2}$, which is of the same size as the model parameters ${\bm{\theta}}$. We believe this issue is of lesser consequences as Fully-Sharded-Data-Paralellism (Zhao et al., 2023, FSDP) can always be used to distribute the optimizer states across compute nodes. In the cases where memory remains an issue, we see two mitigation strategies. The first one consists in setting $\beta_{1}=0$, which removes entirely the need for ${\bm{m}}_{1}$. We show in App. C.1.4 (Fig. 15(b)) that this strategy might work, but at the cost of less stable training. The second mitigation strategy is to apply factorization strategies as in Shazeer & Stern (2018); Zhao et al. (2024).
Hyperparameter sensitivity. While we introduce up to four new hyperparameters: $\alpha$, $\beta_{3}$, $T_{\alpha}$, and $T_{\beta_{3}}$. In practice we always set $T_{\alpha}=T_{\beta_{3}}=T_{\alpha,\beta_{3}}$, and use $T_{\alpha,\beta_{3}}=T$ in most cases. We show in App. C.1.4 that $T_{\alpha}$ and $T_{\beta_{3}}$ should only be large enough to prevent instabilities early during training. While all of our experiments on language modeling use $\beta_{3}=0.9999$, other values such as $0.999$ or even $0.99999$ still can outperform the AdamW baseline (see App. C.1.4, Fig. 12). On vision tasks, the scatter plots in Fig. 6 show all the AdEMAMix models trained for those experiments, highlighting how easy it can be to find good $\alpha$ and $\beta_{3}$ values. Overall, we find the ranges of values of $\alpha,\beta_{3}$ and $T_{\alpha,\beta_{3}}$ providing improvements over AdamW to be wide. See App. C.1.4 for hyperparameter sweeps.
Limitations. AdEMAMix consists in leveraging very old gradients. Therefore, the method is best suited to settings where the number of iterations is important. We report on this effect in App. C.1.5, additionally showing how smaller values of $\beta_{3}$ (e.g. $\beta_{3}=0.999$) can be better for low iterations scenarios. Moreover, retaining gradient information over many thousands steps can pose a problem in domains requiring fast adaptation to a sudden distribution shift, or general cases in which the distribution is non-stationary.
@@ 4.1 Transformer LLM Training
Figure 5: Training time comparison & starting AdEMAMix from AdamW. In (a), we compare the time required to train $110$M and $1.3$B parameter models for $256k$ iterations. The additional EMA renders AdEMAMix slightly slower than AdamW. However, if we were to train AdamW longer to compensate for this gap, we would only train for an additional $2379$ and $5421$ iterations for resp. $110$M and $1.3$B parameter models. Those additional iterations would not be sufficient to close the gap (see Fig. 1). In (b) and (c), we show—for two different model sizes—the effect of switching from AdamW to AdEMAMix during training. AdEMAMix’s additional parameter ${\bm{m}}_{2}$ is initialized to $\mathbf{0}$, no scheduler is used for $\alpha$ or $\beta_{3}$. For both model sizes, we observe the loss increases slightly at first before decreasing and outperforming the baseline. In both cases, the earlier AdEMAMix is used, the better the final loss. See App. C.1.7 for results using schedulers.
@@
Resume from AdamW vs. training AdEMAMix from scratch. So far we trained AdEMAMix models from scratch. We show it is also possible to switch from an AdamW to an AdEMAMix state in the middle of training. When switching to AdEMAMix at step $t_{switch}$, we initialize ${\bm{m}}_{2}^{(t_{switch})}=\mathbf{0}$ and replace $t$ by $t-t_{switch}$ in the scheduler equations—if those are used. However, we find that schedulers are not required when resuming training and report results without them in the main paper (see App. C.1.7 for more details). In Fig 5(b) and Fig.5(c), we show that (i) it is possible to improve upon the baseline when activating AdEMAMix later during training, and (ii) the earlier the switch, the larger the gain, with diminishing returns. This indicates that the improvement of AdEMAMix cannot be attributed solely to early training dynamics, but rather, late training dynamics seem to play an important role. This is further corroborated by the reverse experiment—which switches from AdEMAMix to AdamW mid-training and show a performance degradation (see results in App. C.1.3).
@@ A.2 Pytorch implementation
The following is a code skeleton for our AdEMAMix optimizer in Pytorch (Paszke et al., 2019).
@@ A.3 Optax implementation
The following is a code skeleton for our AdEMAMix optimizer in Optax, an optimization library based on JAX (DeepMind et al., 2020).
@@ C.1.2 More results on forgetting
Figure 9: Comparing forgetting between AdamW and AdEMAMix. In (a) and (b), we follow—over $10k$ iterations—the loss of training batches $B_{\tau}$ for $\tau\in\{30k,40k,\ldots,250k\}$. We measure the loss right before training on $B_{\tau}$, and then monitor this loss over the next $10k$ steps. For instance, the most yellow curve represents the loss on batch $B_{250k}$ that we track for the last $6k$ iterations of training (we train for $256k$ iterations). The loss is normalized such that the loss for $B_{\tau}$ is 00 and the one at time $\tau+50$ is $-1$. This allows us to overlap curves and observe the tendency of the model to forget batches throughout training. We observe that AdEMAMix improves its loss on training batches $B_{\tau}$ over many timesteps, a striking difference compared to AdamW, which has its loss on $B_{\tau}$ increasing faster. Moreover—likely due to the learning rate scheduler—the models forget significantly slower at the end of training. In this later training stage ($\tau>180k$), the difference between the two optimizers is especially pronounced. While this visualization allows to efficiently compare forgetting at different stages of training, the normalization prevents us from comparing the drop in loss between AdEMAMix and AdamW.
@@ C.1.4 Hyperparameter sensitivity
Figure 13: Sensitivity of AdEMAMix to $T_{\alpha}$ and $T_{\beta_{3}}$. For all the AdEMAMix experiments in this figure, we used $\beta_{1}=0.9,\beta_{2}=0.999,\beta_{3}=0.9999,\alpha=10$. In (a), we test the sensitivity of $110$M parameter models to the values of $T_{\alpha}$ and $T_{\beta_{3}}$. We tied the two values as in all the experiments in the main paper: $T_{\alpha}=T_{\beta_{3}}=T_{\alpha,\beta_{3}}$. We observe that when not enough warmup steps are used, the iterates either diverge or converge to slightly worse loss values. In (b) we set $T_{\beta_{3}}=256k$ and only vary $T_{\alpha}$. We observe that $T_{\alpha}$ needs to be large enough to avoid divergence. This demonstrates the necessity of using a warmup on $\alpha$. This being said, the necessity of increasing $\alpha$ linearly depends on the value of $\alpha$. Using a smaller $\alpha$ value can work with $T_{\alpha}=0$, but using an $\alpha$-scheduler alleviates any instability issue and increases the range of optimal $\alpha$ values (see Fig. 14). Moreover, in other experiments, we observe how not using schedulers yields detrimental large updates in the early training phase, see App. C.1.7, Fig. 22.

### v2

@@ 1 Introduction
The use of momentum in optimization is grounded in the varying nature of gradients. As local linear approximations of the loss landscape, their information can quickly become outdated as the optimization process progresses (Pascanu et al., 2013). For this reason, practitioners typically employ moderate momentum values (i.e. $\beta\approx 0.8$ or $0.9$), effectively creating a moving average of recent gradients while discarding older information. Selecting larger $\beta$ values seems counter-intuitive, as it would suggest that older gradients maintain their relevance over extended periods of training. While it is tempting to see the use of small $\beta$s as a confirmation of the limited temporal relevance of gradients, our work reveals instead that older gradients can efficiently be used. When we increase $\beta$, we decrease the relative importance of recent gradients, and the iterate now fails to respond to local changes in the loss landscape. We observe that a single EMA cannot both give a significant weight to recent gradients, and give a non-negligible weight to older gradients (see Fig. 3(a)). However, a linear combination between a “fast-changing” (e.g. $\beta=0.9$ or $\beta_{1}=0$) and a “slow-changing” (e.g. $\beta=0.9999$) EMA allows the iterate to beneficiate from (i) the great speedup provided by the larger (slow-changing) momentum, while (ii) still being reactive to small changes in the loss landscape (fast-changing). More precisely, we find the following statement to convey an important intuition behind this approach: {adjustwidth}2em2em While changing the direction of the slow momentum is difficult, any adjustment orthogonal to that direction is easy—which favors fast progress in sinuous canyon-like landscapes. A toy illustration of this can be seen in Fig. 2. Based on this idea, we propose AdEMAMix (Adaptive EMA Mixture), a novel Adam based optimizer which successfully leverages very old gradients to reach better solutions.
@@
  • We show AdEMAMix forgets the training data slower when compared to Adam (see Fig. 4).
@@ 2 Related Work
Work incorporating an additional momentum term. Lee et al. (2024) introduce Grokfast, which uses a pre-filtering step on the gradient to amplify the low frequencies and help solve groking. When combined with Adam, it effectively applies the Adam’s EMAs on top of another gradient averaging method (e.g. EMA for Grokfast-EMA). Somewhat similarly, Chen et al. (2023b) refer to the Double EMA (DEMA) used in some finance applications (Mulloy, 1994) as one motivation for their AdMeta optimizer. Our motivation behind AdEMAMix is to combine both a high sensitivity to the recent gradients as well as incorporating very distant gradient, in this respect, using nested EMAs is not the right candidate as it reduces the influence of recent gradients. A more detailed review of AdMeta and DEMA is in App. C.3.1. Most relevant to us, Lucas et al. (2019, AggMo) also observe that using a combination of EMAs can enable the use of larger $\beta$s, and incorporates a sum of $K$ momentum terms into GD They show their approach reaches similar performances as baseline optimizers, with a faster convergence. In contrast, we modify Adam, and introduce schedulers that are critical to reaching good performances at larger scales. This allows us to outperform Adam by a significant margin. In App. C.3.3, we notice no further improvement by adding more momentum terms. Finally, Szegedy et al. (2024) propose a general framework to derive and study optimizers with linear combinations of memory vectors—which encompasses EMA mixtures.
@@ 3 Our method: AdEMAMix
Figure 3: Limitation of EMAs, constant $\eta$-scheduler, & Mamba results. In (a), we plot the weights $w_{t}$—for each past gradient ${\bm{g}}^{(t)}$—given by different EMAs after $10k$ steps. For a given $\beta$, $\text{EMA}(\beta,{\bm{g}}^{(0)},\ldots,{\bm{g}}^{(T)})=\sum_{i=0}^{T}\beta^{i}(1-\beta){\bm{g}}^{(T-i)}$, which decays the contribution of past gradients exponentially. A small $\beta$ (e.g. $0.9$) will give a high weight to the immediate past and negligible weights to older timesteps. In contrast, a high $\beta$ (e.g. $0.9999$) is giving a relatively uniform, yet non-negligible weight to all gradients. No $\beta$ value can simultaneously give a high weight to the immediate past and a non-negligible weight to very old timesteps. In (b), we train multiple $1.3$B language models using $3k$ steps of warmup and then a constant learning rate $\eta=10^{-4}$. This allows us to observe the gap between AdamW and AdEMAMix without the cosine decay as a confounder. We still use schedulers for $\alpha$ and $\beta_{3}$ with $T_{\alpha,\beta_{3}}=500k$, $\alpha=5$. Similar to Zhai et al. (2022); Hu et al. (2024); Hägele et al. (2024), we decay the learning rate linearly at $t=1$M and $t=1.3$M. The loss-gap between AdamW and AdEMAMix increases at first, and then remains constant. AdEMAMix still outperforms AdamW after decaying the learning rate. See App.C.1.9 to see the impact of the linear decay duration. In (c), we train $168$M parameter Mamba models, showing how AdEMAMix’s performances can generalize outside of the Transformer architecture.
@@
Tackling early training instabilities. Early training instabilities are commonplace when training deep models, and empirically motivated the introduction of methods such as learning rate warmup and gradient clipping. Gilmer et al. (2022) show how the use of learning rate warmup can be justified from a curvature perspective; allowing the iterates to move to parts of the optimization landscape where larger learning rates are stable. While we use learning rate warmup in all our experiments, we still noticed AdEMAMix models using a large $\beta_{3}$ would diverge early. This, despite not using bias correction over ${\bm{m}}_{2}$, which lets the momentum buffer fill itself slowly. Those failed runs are characterized by updates of large magnitudes in the early phase of training (see App. C.1.9, Fig. 27). For this reason, we progressively increase the values of $\beta_{3}$ and $\alpha$ using schedulers. For $\alpha$ we use a linear scheduler. A linear scheduler for $\beta_{3}$ would be ill-fitted as the same increment of $\beta_{3}$ has a different impact for different values of $\beta_{3}$. For instance, observe that an increase of $\beta$ of $\delta_{\beta}=0.0001$ barely increases the $t_{half}$ for $\beta=0.9$, while $0.999\rightarrow 0.999+\delta_{\beta}$ increases the $t_{half}$ of $77$. For this reason, we design the $\beta_{3}$ scheduler to increase $t_{half}$ linearly (see App. A.1 for a derivation of that scheduler). The two schedulers are summarized below:
@@
Using $\beta_{1}=0$ to save memory. By setting $\beta_{1}=0$, we can save on memory by replacing $\hat{{\bm{m}}}_{1}^{(t)}$ by ${\bm{g}}^{(t)}$. In this case, the memory cost of AdEMAMix is the same as for AdamW. We show in App. C.1.4 (Fig. 15(b)) and App. C.1.7 that using $\beta_{1}=0$ often works, at the cost of potentially less stable training.
Computational overheads. Adding an additional EMA requires additional memory and compute. The added compute is negligible in comparison to what is required for the forward-backward, and has little impact over the total runtime as shown in Fig. 5(a). Moreover—when considering larger distributed setups—it is worth noting that AdEMAMix is not increasing communication (gradient reduction) over Adam. Therefore, we expect the runtime overhead of AdEMAMix to shrink in those settings, as data movements occupy a larger fraction of the total runtime. A more significant overhead is in terms of memory when $\beta_{1}\neq 0$, as AdEMAMix requires to allocate both ${\bm{m}}_{1}$ and ${\bm{m}}_{2}$, which are of the same size as the model parameters ${\bm{\theta}}$. We believe this issue is of lesser consequences as Fully-Sharded-Data-Paralellism (Zhao et al., 2023, FSDP) can always be used to distribute the optimizer states across compute nodes. In the cases where memory remains an issue, one mitigation strategy could be to apply factorization strategies as in Shazeer & Stern (2018); Zhao et al. (2024).
Hyperparameter sensitivity. While we introduce up to four new hyperparameters: $\alpha$, $\beta_{3}$, $T_{\alpha}$, and $T_{\beta_{3}}$, in practice we always set $T_{\alpha}=T_{\beta_{3}}=T_{\alpha,\beta_{3}}$, and use $T_{\alpha,\beta_{3}}=T$ in most cases. We show in App. C.1.4 that $T_{\alpha}$ and $T_{\beta_{3}}$ should only be large enough to prevent instabilities early during training. While all of our experiments on language modeling use $\beta_{3}=0.9999$, other values such as $0.999$ or even $0.99999$ still can outperform the AdamW baseline (see App. C.1.4, Fig. 12). On vision tasks, the scatter plots in Fig. 6 show all the AdEMAMix models trained for those experiments, highlighting how easy it can be to find good $\alpha$ and $\beta_{3}$ values. Overall, we find the ranges of values of $\alpha,\beta_{3}$ and $T_{\alpha,\beta_{3}}$ providing improvements over AdamW to be wide. See App. C.1.4 for hyperparameter sweeps.
Limitations. AdEMAMix consists in leveraging very old gradients. Therefore, the method is best suited to settings where the number of iterations is important. We report on this effect in App. C.1.5, additionally showing how smaller values of $\beta_{3}$ (e.g. $\beta_{3}=0.999$) can be better for low iterations scenarios. Moreover, retaining gradient information over many thousands steps can pose a problem in domains requiring fast adaptation to a sudden distribution shift.
@@ 4.1 Transformer LLM Training
Figure 5: Training time comparison & starting AdEMAMix from AdamW. In (a), we compare the time required to train $110$M and $1.3$B parameter models for $256k$ iterations. The additional EMA renders AdEMAMix slightly slower than AdamW. However, if we were to train AdamW longer to compensate for this gap, we would only train for an additional $2379$ and $5421$ iterations for resp. $110$M and $1.3$B parameter models. Those additional iterations would not be sufficient to close the gap (see Fig. 1). In (b) and (c), we show—for two different model sizes—the effect of switching from AdamW to AdEMAMix during training. AdEMAMix’s additional parameter ${\bm{m}}_{2}$ is initialized to $\mathbf{0}$, no scheduler is used for $\alpha$ or $\beta_{3}$. For both model sizes, we observe the loss increases slightly at first before decreasing and outperforming the baseline. In both cases, the earlier AdEMAMix is used, the better the final loss. See App. C.1.9 for results using schedulers.
@@
Resume from AdamW vs. training AdEMAMix from scratch. So far we trained AdEMAMix models from scratch. We show it is also possible to switch from an AdamW to an AdEMAMix state in the middle of training. When switching to AdEMAMix at step $t_{switch}$, we initialize ${\bm{m}}_{2}^{(t_{switch})}=\mathbf{0}$ and replace $t$ by $t-t_{switch}$ in the scheduler equations—if those are used. However, we find that schedulers are not required when resuming training and report results without them in the main paper (see App. C.1.9 for more details). In Fig 5(b) and Fig.5(c), we show that (i) it is possible to improve upon the baseline when activating AdEMAMix later during training, and (ii) the earlier the switch, the larger the gain, with diminishing returns. This indicates that the improvement of AdEMAMix cannot be attributed solely to early training dynamics, but rather, late training dynamics seem to play an important role. This is further corroborated by the reverse experiment—which switches from AdEMAMix to AdamW mid-training and show a performance degradation (see results in App. C.1.3).
@@ A.2 Pytorch implementation
The following is a code skeleton for our AdEMAMix optimizer in Pytorch (Paszke et al., 2019). The full implementation of AdEMAMix in Pytorch is provided in the following repository: https://github.com/apple/ml-ademamix
@@ A.3 Optax implementation
The following is a code skeleton for our AdEMAMix optimizer in Optax, an optimization library based on JAX (DeepMind et al., 2020). The full implementation of AdEMAMix in Jax is provided in the following repository: https://github.com/apple/ml-ademamix
@@ C.1.2 More results on forgetting
Figure 9: Comparing forgetting between AdamW and AdEMAMix. In (a) and (b), we follow—over $10k$ iterations—the loss of training batches $B_{\tau}$ for $\tau\in\{30k,40k,\ldots,250k\}$. We measure the loss right before training on $B_{\tau}$, and then monitor this loss over the next $10k$ steps. For instance, the most yellow curve represents the loss on batch $B_{250k}$ that we track for the last $6k$ iterations of training (we train for $256k$ iterations). The loss is normalized such that the loss for $B_{\tau}$ is 00 and the one at time $\tau+50$ is $-1$. This allows us to overlap curves and observe the tendency of the model to forget batches throughout training. We observe that AdEMAMix improves its loss on training batches $B_{\tau}$ over many timesteps, a striking difference compared to AdamW, which has its loss on $B_{\tau}$ increasing faster. Moreover—likely due to the learning rate scheduler—the models forget significantly slower at the end of training. In this later training stage ($\tau>180k$), the difference between the two optimizers is especially pronounced. While this visualization allows to efficiently compare forgetting at different stages of training, the normalization prevents us from comparing the drop in loss between AdEMAMix and AdamW. To compare loss values, see Fig. 4.
@@ C.1.4 Hyperparameter sensitivity
Figure 13: Sensitivity of AdEMAMix to $T_{\alpha}$ and $T_{\beta_{3}}$. For all the AdEMAMix experiments in this figure, we used $\beta_{1}=0.9,\beta_{2}=0.999,\beta_{3}=0.9999,\alpha=10$. In (a), we test the sensitivity of $110$M parameter models to the values of $T_{\alpha}$ and $T_{\beta_{3}}$. We tied the two values as in all the experiments in the main paper: $T_{\alpha}=T_{\beta_{3}}=T_{\alpha,\beta_{3}}$. We observe that when not enough warmup steps are used, the iterates either diverge or converge to slightly worse loss values. In (b) we set $T_{\beta_{3}}=256k$ and only vary $T_{\alpha}$. We observe that $T_{\alpha}$ needs to be large enough to avoid divergence. This demonstrates the necessity of using a warmup on $\alpha$. This being said, the necessity of increasing $\alpha$ linearly depends on the value of $\alpha$. Using a smaller $\alpha$ value can work with $T_{\alpha}=0$, but using an $\alpha$-scheduler alleviates any instability issue and increases the range of optimal $\alpha$ values (see Fig. 14). Moreover, in other experiments, we observe how not using schedulers yields detrimental large updates in the early training phase, see App. C.1.9, Fig. 27.